### v2

# Disorder-induced proximate quantum spin ice phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$

Abstract
We report a comprehensive bulk characterization and neutron scattering investigation of single-crystalline $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, a magnetic pyrochlore synthesized via a flux-growth method. Unpolarized neutron diffuse scattering reveals the emergence of spin-ice correlations below $T\sim 1$ K, evidenced by the development of anisotropic pinch-point features that are consistent with quantum-spin-ice (QSI) behavior. A.C. susceptibility measurements indicate a progressive slowing of spin dynamics in this regime, culminating in complete spin freezing below $T_{f}\approx 0.15$ K. Inelastic neutron scattering at $T=0.5$ K reveals a broad spectrum of quasi-elastic magnetic excitations, with intensity in the low-energy range $[0,0.2]$ meV significantly suppressed below $T_{f}$. Meanwhile, an incipient (100)-type magnetic order begins to nucleate, and a gapped excitation centered at $\hbar\omega=0.23$ meV persists. We further identify two distinct dynamical timescales above $T_{f}$, a slow component $\tau_{\mathrm{slow}}\sim 10^{-5}$ s and a fast component $\tau_{\mathrm{fast}}\sim 10^{-10}$ s, in quantitative agreement with theoretical predictions for QSI systems. Taken together, these results indicate that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  enters a disorder-induced spin-frozen phase below $T_{f}$, lying in close proximity to a $U(1)$ quantum spin liquid.
I Introduction
The discovery of classical spin ice (CSI) phases in rare-earth pyrochlore compounds marks a major success in the search for exotic magnetic phases with potential applications in quantum computing [1] and magnetic storage technologies [2]. The ground state of CSI is characterized by the “two-in, two-out” ice rules for semi-classical magnetic moments residing on corner-sharing tetrahedra and pointing along the local $\langle 111\rangle$ directions. In the first established CSI compounds, $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$ [3] and $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$ [4], this ground state results from dominant Ising anisotropy and the self-screening dipolar interaction, which together generate ferromagnetic nearest-neighbor interactions. The CSI phase possesses a massive ground state degeneracy analogous to the degeneracy of proton order in water ice. It also exhibits dipolar-like spin-spin correlations that manifest as pinch-point features in magnetic neutron scattering patterns [3].
The success of CSI has inspired the theoretical proposal of quantum spin ice (QSI) [5], where anisotropic exchange interactions between spins enable quantum tunneling between different ice-rule-obeying configurations, thereby stabilizing a U(1) quantum spin liquid (QSL) phase with emergent electrodynamic properties. The QSI phase supports both gapless, photon-like magnetic excitations and gapped topological defects known as “magnetic monopoles” (or “visons,” as referred to in this work) [6, 7]. These monopoles represent topological defects of the emergent vector gauge potential in the U(1) QSL and have no classical analogue.
Figure 1: A qualitative phase diagram of the antiferromagnetic Ising model with random transverse field as defined in Eq. 1 with only nearest-neighbor interaction $J_{1}>0$. The parameter $\bar{h}$ denotes the spatial average of the site-dependent transverse fields $h_{i}$, while $\delta h$ characterizes the width of their spatial distribution. The red solid line marks a first-order Higgs transition, corresponding to global spinon condensation. The red dashed line indicates a crossover regime of subregional spinon condensation. The blue solid line represents a confinement transition of the emergent U(1) gauge field, associated with the condensation of visons and the onset of a glassy phase with frozen magnetic moments.
Rare earth pyrochlores have featured prominently in the search for QSI candidates. Notable examples include $\text{Yb}_{2}\text{Ti}_{2}\text{O}_{7}$ [8, 9], the Ce-based compounds $\text{Ce}_{2}\text{Zr}_{2}\text{O}_{7}$  [10, 11, 12, 13, 14, 15], $\text{Ce}_{2}\text{Sn}_{2}\text{O}_{7}$ [16, 17, 18], and $\text{Ce}_{2}\text{Hf}_{2}\text{O}_{7}$ [19, 20], and the Pr-based compounds $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  [21, 22, 23, 24] and $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  [25, 26, 27]. The QSI candidacy of $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  was initially supported by the observation of narrow pinch points in the elastic magnetic scattering, and over 90% of the spectral weight in an inelastic continuum devoid of pinch-point features [21, 23]. In contrast, the reported single crystals of its sister compound $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  exhibit predominantly elastic magnetic scattering, accompanied by a gapped excitation near $\hbar\omega=0.2$ meV [25, 27]. Subsequent studies [22, 24] have revealed that disorder plays a prominent role in determining the ultra-low-temperature magnetic behavior of Pr-based pyrochlore magnets. In these systems, the non-Kramers ground state doublets of Pr3+ ions [21, 26, 28] can be split by any perturbation that breaks the local $D_{3d}$ symmetry. This effect can be modeled as a site-dependent transverse field $h_{i}$ acting on each Pr3+ ion, described by the Hamiltonian
$\begin{split}\mathcal{H}=\sum_{\langle ij\rangle,n}J_{n}\sigma_{i}^{z}\sigma_{j}^{z}-\sum_{i}h_{i}\sigma_{i}^{x}.\end{split}$ (1)
The longitudinal pseudospin operator $\sigma_{i}^{z}$ is related to the local magnetic moment through $\bm{m}_{i}=\mu_{\mathrm{eff}}\sigma_{i}^{z}\hat{\bm{z}}_{i}$, where $\mu_{\mathrm{eff}}$ denotes the effective moment of the ground-state doublet, and $\hat{\bm{z}}_{i}$ is the unit vector along the local $\langle 111\rangle$ direction (following the convention of Ref. [8]). The transverse pseudospin components $\sigma_{i}^{x}$ and $\sigma_{i}^{y}$ represent quadrupolar degrees of freedom [23] and do not contribute to the magnetic dipole moment [29, 30]. Nearest neighbor ($n=1$) antiferromagnetic interactions $J_{1}>0$ between pseudospins (equivalent to ferromagnetic interactions between magnetic moments) stabilize the “two-in, two-out” spin-ice state. In Eqn. 1, we have omitted the nearest-neighbor transverse interaction that is symmetrically allowed in Pr-based systems [23, 24]. Throughout this paper, we assume that a finite pairwise interaction between transverse pseudospin components may also be present, but focus primarily on the influence of $J_{1}$ and $h_{i}$.
The interplay of spin-ice correlations and disorder-induced transverse fields can yield a rich space of QSL and confined phases [31, 29, 30, 32]. A qualitative phase diagram, based on Refs. [29, 30, 32], is presented in Fig. 1, where $\bar{h}$ denotes the spatial average of the site-dependent transverse field $h_{i}$, while $\delta h$ characterizes the width of its spatial distribution. Starting from the clean limit, a small uniform transverse field $h_{i}$ enhances quantum fluctuations and can help stabilize the QSI phase [29] by generating a “random ring-exchange” term that facilitates quantum tunneling within the spin-ice manifold and enables the emergence of gapped topological excitations (visons). Two instabilities can drive the system out of the QSI phase, corresponding to the condensation of either spinons (spin flip defects) or visons. When $\bar{h}\gtrsim h_{c}\approx 0.6J_{1}$, spinons condense and the system transitions into a paramagnetic phase consisting of site-localized singlets [31, 30].In contrast, a finite $\delta h$ introduces spatial inhomogeneity in the ring-exchange amplitudes and enhances vison dispersion. Beyond a critical threshold $\delta h_{c}$, the system undergoes a confinement transition where the visons condense [29, 30]. This favours ordering with a (100) propagation vector [6]; however, long-range (100) order is frustrated by structural disorder and the system instead enters a glassy phase with frozen magnetic moments. For QSI candidates, the critical disorder variance $\delta h_{c}$ at vanishing average field $\bar{h}\rightarrow 0$ is expected to remain finite due to quantum fluctuations arising from intrinsic transverse interactions. In systems with finite $\delta h$, an additional “Griffiths phase” [33, 29] may emerge, in which rare spatial regions locally satisfy $\bar{h}>h_{c}$ and enter the paramagnetic phase, while the bulk of the system either remains in a fluctuating QSI regime (“Griffiths QSI phase”) or transitions into a frozen state due to disorder-induced confinement (“Griffiths spin-frozen phase”).
To realize a QSI state experimentally, it is essential to characterize the phases proximate to the QSI regime and to establish where candidate materials lie within the phase diagram. Previous studies [22, 30] place $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  deep within the paramagnetic phase. In contrast, $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  samples—which exhibit significantly less inelastic scattering [27] and clear signatures of spin freezing [25]—have not been analyzed within the framework shown in Fig. 1. Among the Pr-based pyrochlores, $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  has long been considered a promising QSI candidate: results on powder samples suggest that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  is a promising QSI candidate, since spins remain fluctuating down to at least 0.2 K, with no evidence of long-range magnetic ordering [34, 35]. However, large single crystals of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  suitable for inelastic neutron scattering have not previously been available.
In this work, we present the first experimental study of flux-grown single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, combining comprehensive neutron scattering and thermodynamic measurements. Our key findings are as follows. First, the magnetic diffuse scattering observed within the spin-ice regime at $T\sim 0.3$ K cannot be accounted for by the classical nearest-neighbor spin-ice model; instead, either quantum fluctuations or further-neighbor interactions must be invoked. Second, magnetic susceptibility and inelastic neutron scattering measurements reveal two well-separated dynamical timescales, consistent with theoretical predictions for QSI states [36] and with previous powder-sample results [35], thereby lending strong support to the QSI scenario. Finally, below $T_{f}\approx 0.15$ K, our flux-grown single crystals enter a disorder-induced spin-frozen state with incipient $(100)$ magnetic order, positioned in close proximity to the QSI phase. This behavior contrasts sharply with earlier reports on floating-zone $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ samples, which remain firmly in the paramagnetic regime [22, 30].
II Methods
Single crystals of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  were grown via a flux method with a 1:1.2 mass ratio of dried Na2B4O7:NaF as the flux. This was mixed with Pr6O11, SnO2, and SnO to yield a nominal 1:5 mass ratio of Pr2Sn2O7 to flux. The mixture was ground and loaded into a 20 mL platinum crucible with a lightly fitting platinum lid, heated to 1000C at 200C/hr, soaked for 6 hours, and cooled to 500C at 1–1.5C/hr. Approximately 50% of the flux mass was lost due to volatilization. Residual flux was removed by dissolving in boiling water. The resulting crystals are well-faceted dark red octahedra, typically 1–20 mg in mass and millimeter-sized. A single-crystal specimen with a mass of approximately 1 mg was used for heat capacity and a.c. susceptibility measurements using a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS), covering both the conventional 4He and dilution refrigerator (DR) temperature ranges. For single-crystal neutron scattering experiments, a total of 175 single-crystal pieces were co-aligned with the $(HHL)$ scattering plane oriented horizontally, yielding a combined mass of approximately 1.27 g and a mosaicity of $\sim 5^{\circ}$ full width at half maximum (FWHM). Additional details about the samples are provided in Appendix C and Fig. A2.
Magnetic diffuse scattering measurements were performed on the wide-angle single-crystal diffractometer WAND2 of the High Flux Isotope Reactor (HFIR) at Oak Ridge National Laboratory (ORNL), using unpolarized neutrons with an incident wavelength of $\lambda=1.5$$\text{\AA }$. Two primary data sets were collected at $T=0.3$ K and $T=1$ K, with background subtracted using high-temperature measurements taken at $T=10$ K and $T=20$ K, respectively. The use of two distinct background data sets compensates for a minor unintended change in background conditions during the experiment. We note that for $T\geq 10$ K, no observable spin-ice correlations are expected to form in reciprocal space (see Fig. 4(ii)).
Inelastic neutron scattering measurements were performed on the Cold Neutron Chopper Spectrometer (CNCS) at the Spallation Neutron Source (SNS), ORNL. Data were collected at three temperatures: $T=0.05$ K, $T=0.5$ K, and $T=10$ K. For each temperature, measurements were carried out using two incident neutron energies, $E_{i}=0.99$ meV and $E_{i}=3.32$ meV, both using the high-flux chopper configurations. These settings provided energy resolutions of approximately 0.02 meV and 0.1 meV at the elastic line, respectively. For both the WAND2 and CNCS experiments, the sample was rotated about the vertical $1\bar{1}0]$ axis over a total range of $180^{\circ}$, using step sizes of $0.1^{\circ}$ and $1^{\circ}$, respectively. The data were reduced and processed using the Mantid software suite [37].
To assess the nature and extent of structural disorder in our sample, powder neutron-diffraction measurements were performed on the Nanoscale-Ordered Materials Diffractometer (NOMAD) at SNS, ORNL. Data were collected from crushed single-crystal specimens with a total mass of $\sim$0.5 g at $T=100$ K and $T=300$ K. Rietveld refinements of the diffraction patterns and pair-distribution function (PDF) analysis were carried out using the TOPAS  [38] and PDFgui [39] software packages, respectively.
III Results
Figure 2: Pair distribution function (PDF) analysis of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  obtained from total neutron scattering data collected on the Nanoscale-Ordered Materials Diffractometer (NOMAD) at $T=100$ K. Shown are the experimental data (black), the refined fit (red), and the difference curve (blue).
Figure 3: Magnetic diffuse scattering of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  from experiment (a,e) and simulations (b–d, f–h). Panels (a) and (e) show experimental data at $T=0.3$ K and $T=1$ K, respectively, with high-temperature backgrounds ($T=10$ K and $T=20$ K) subtracted. Data are integrated along $K\bar{K}0]$ over $K\in[-0.1,0.1]$ rlu. Simulations include: (b,c,f,g) the quantum spin ice model from Ref. [40], with (b,f) assuming the quantum limit and perturbative transverse interactions with uniform ring exchange $K^{\mathrm{eff}}=0.075(1)$ K; (c,g) the classical spin ice model with nearest-neighbor interaction $J_{1}$ only; and (d, h) best-fit results for dipolar spin ice (DSI) model with the calculated dipolar interaction $D_{\mathrm{dip}}\approx 0.012$ K (following the convention in Ref.[41]), the exchange interaction $J_{1}^{\mathrm{ex}}=0.25(1)$ K, and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K (see Eq. 1 and Appendix. D for the conventions).
III.1 Crystal-structure refinement
Rietveld refinement of neutron powder diffraction data collected on a crushed single crystal reveals no measurable signatures of site disorder for models typical of pyrochlore materials, including substitution of Pr4+ on the Pr site, corresponding to Pr2Sn2O7+y; antisite cation disorder, modeled as Pr2(Sn2-xPrx)O7+y; or Frenkel-type oxygen disorder. These results place an upper limit of $\approx 1\%$ site disorder for the above models. The only disorder model consistent with neutron powder-diffraction data invokes $\sim$5% of Pr-site vacancies; however, this value is uncertain due to large parameter correlations, and refinements to single-crystal X-ray diffraction data indicate a smaller value ($\sim$1%) of Pr-site vacancies that is likely more robust (see Appendix E for details).
Pair distribution function (PDF) analysis provides a complementary probe of local disorder. The neutron PDF pattern collected at 100 K is shown in Fig. 2, together with a fit using the stoichiometric model over the first 10 Å of the data. The fit quality is good ($R_{w}=7.71\%$), confirming the adequacy of the model. Nevertheless, small but systematic features in the fit residuals indicate subtle deviations from the nominal structure. Although an unambiguous identification of the origin of this local disorder lies beyond the scope of the present work, our analysis suggests several possibilities. In particular, we found it necessary to refine the symmetry-allowed anisotropic terms in the atomic displacement parameter (ADP) tensor to achieve a good fit for many peaks, most notably the Pr–O and O–O correlations at 2.587 Å and 2.727 Å, respectively, pointing to a relatively broad and anisotropic distribution of these bond lengths. Interestingly, inclusion of anisotropic ADPs worsened the fit for the metal–metal peak at the nearest-neighbor distance of 3.746 Å compared to the isotropic case, although the overall agreement remained substantially better with anisotropic ADPs ($R_{w}=9.33\%$ for the isotropic model). This suggests that the local disorder, whatever its microscopic origin, is primarily accommodated by shifts of the oxygen atoms, leading to slightly distorted PrO8 and/or SnO8 polyhedra while preserving the metal-atom positions expected from the average structure. Finally, we note that the refined lattice parameter exhibits a weak dependence on the maximum $r$-range of the fit (10–20 Å), consistent with the presence of local strains in the sample. Refinment results of both Rietveld refinment and PDF results are provided in Table. 1a 2d and Fig. A6 in Appendix. E.
III.2 Magnetic diffuse scattering
The single-crystal magnetic diffuse scattering measured at $T=0.3$ K and $T=1$ K is shown in Fig. 3(a) and 3(e), respectively. At $T=1$ K, the $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ sample already exhibits a characteristic “pinch-point” pattern, indicative of emerging spin-ice correlations. At $T=0.3$ K [Fig. 3(a)], a more asymmetric pattern develops: compared to the nearest-neighbor CSI simulation [40] [Fig. 3(c)], we observe suppression of the pinch points, anisotropic intensity along the $[001]$ and $[111]$ directions, and broader scattering near the (222) position. Hence, our $T=0.3$ K data cannot be well described using the nearest-neighbor CSI model. To investigate the origin of the observed wavevector dependence, we performed fits using two distinct models. First, we examined the QSI model of Ref. [40] characterized by a uniform ring-exchange interaction with fitted value $K^{\mathrm{eff}}=0.075(1)$ K (defined in Ref. [40]).
Second, a classical model was implemented [41] incorporating fixed dipolar interactions $D_{\mathrm{dip}}\approx 0.012$ K (following the convention in Ref.[41]) and fitted exchange couplings $J_{1}^{\mathrm{ex}}=0.25(1)$ K, and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K (see Appendix D and Figs. A4 and A5 for details and conventions). Both the quantum model [Fig. 3(b,f)] and the classical model with next-nearest neighbor exchange interactions [Fig. 3(d,h)] provide an excellent description of our data.
On the basis of equal-time magnetic correlations alone, it is therefore not possible to disentangle the role of quantum dynamics from that of further-neighbor interactions. To resolve this ambiguity, we turn to the dynamical properties. Theory predicts [36] that a QSI phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ should exhibit two well-separated timescales: a slow component, $\tau_{\mathrm{slow}}$, detectable by a.c. susceptibility, and a fast component, $\tau_{\mathrm{fast}}$, observable by inelastic neutron scattering. We therefore collected single-crystal data using each of these probes.
Figure 4: Bulk characterization of flux-grown single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$. (a) Temperature dependence of the real part of the a.c. susceptibility $\chi^{\prime}$ measured along the $[111]$ direction at various frequencies and zero dc field. Data are normalized to the dc susceptibility between 1.8–4 K (black curve). Inset (i): Temperature dependence of the “slow” spin relaxation times [36] of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ extracted from (a), fitted using a damped harmonic oscillator model [42, 35]. (b) Imaginary part $\chi^{\prime\prime}$ of the a.c. susceptibility under the same conditions. (c) Arrhenius fit ($f=f_{0}\exp(-E_{a}/T_{f})$) of the $\chi^{\prime\prime}$ peak frequency, yielding an activation energy $E_{a}=2.7(7)$ K. Frequencies used in (a,b) are indicated in the color bar. (d) Heat capacity as a function of temperature after subtracting the lattice contribution estimated from non-magnetic La2Sn2O7, compared with two models for the nuclear Schottky anomaly: (1) the Kimura model [21], using $\mu_{\mathrm{hyp}}^{\mathrm{Pr}}=1.9~\mathrm{\upmu_{B}}$; and (2) the Gopal model [43, 35], assuming evenly spaced levels with $\epsilon_{r}=0.011$ meV. Inset (ii): Magnetic entropy obtained by subtracting the Gopal model from (d) and integrating $C_{m}/T$, recovering the full Pauling entropy near 10 K.
III.3 Bulk characterization
As shown in Fig. 4(a,b), a.c. susceptibility measurements on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  with zero d.c. field reveal a gradual slowing down of spin dynamics below 1 K. This is manifested in the frequency-dependent behavior of the real part, $\chi^{\prime}$, whose peak shifts to higher temperatures with increasing frequency. A complete spin-freezing transition occurs at $T_{f}\approx 0.15$ K, where $\chi^{\prime}$ drops nearly to zero. By applying a damped-harmonic oscillator model [42, 35] to $\chi^{\prime}$, we estimate that $\tau_{\mathrm{slow}}$ ranges from $10^{-6}$ s to $10^{-2}$ s below 1 K [Fig. 4(i), Fig. A3]. The frequency dependence of the freezing temperature $T_{f}$, determined from the position of the cusp in the imaginary part $\chi^{\prime\prime}$ shown in Fig. 4(b), follows Arrhenius behavior with activation energy $E_{a}=2.7(7)$ K [Fig. 4(c)], consistent with previous results on powder samples [35].
The zero-field heat capacity $C_{p}$ is shown in Fig. 4(d) and exhibits a pronounced nuclear Schottky anomaly that approaches the theoretical maximum value of $\sim 7~\mathrm{J}/(\mathrm{mol}_{\mathrm{Pr^{3+}}}\cdot\mathrm{K})$, corresponding to fully frozen Pr moments [21, 44] at base temperature. The recovered zero-point entropy, obtained by subtracting the hyperfine contribution (see Appendix B), approaches the full Pauling value near 10 K. Taken together, these results suggest that while the system enters a fully spin-frozen state at base temperature, a quantum spin-ice (QSI) phase may emerge above the spin-freezing transition, characterized by a “slow” timescale of $\tau_{\mathrm{slow}}\sim 10^{-5}$ s [36].
Figure 5: Neutron spectroscopy of single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$. (a–c) Powder-averaged scattering intensity as a function of energy transfer $\hbar\omega$ at $T=0.05$, 0.5, and 10 K, measured with $E_{i}=0.99$ meV (elastic resolution FWHM $\approx 0.02$ meV). Data are integrated along $K\bar{K}0]$ with $K\in[-0.1,0.1]$ rlu and over $Q\in[0.05,1.3]~\mathrm{\AA }^{-1}$ within the horizontal $(HHL)$ plane. For visual comparison, the $T=0.5$ K data from (b) are also shown in (a). Insets (i,$i$,iii) display the corresponding elastic intensities. The best-fit inelastic contribution using Eq. 3, with $E_{\mathrm{gap}}=0.23(1)$ meV and $\Gamma=0.11(2)$ meV, is shown in (a). Quasielastic fits based on Eq. 2 are shown in (b) and (c), yielding $\Gamma=0.062(3)$ meV at $T=0.5$ K and $\Gamma=0.10(1)$ meV at $T=10$ K. (d-j) Constant-energy $\bm{Q}$-maps at $T=0.5$ K (d, f, h) and $T=0.05$ K (e, g, j), measured with $E_{i}=3.32$ meV (elastic resolution FWHM $\approx 0.1$ meV) after subtraction of the $T=10$ K background. Data are integrated along $K\bar{K}0]$ with $K\in[-0.15,0.15]$ rlu and over the energy ranges $\hbar\omega\in[-0.2,0.2]$ meV (d, e), $[0.2,0.6]$ meV (f, g), and $[0.6,1.0]$ meV (h, j). No symmetrization is applied. Color scales are centered on the mean intensity of each panel. Energy-gain and energy-loss contributions are combined in panels (f,g,h,j); the two channels are displayed separately in Fig. A12.
III.4 Neutron spectroscopy
Neutron spectroscopy provides crucial insights into the spin dynamics above and below $T_{f}$. We first consider the energy dependence of the scattering using $Q$-averaged energy cuts from high-resolution $E_{i}=0.99$ meV measurements, as shown in Fig. 5(a–c), where the data are integrated over the $Q$-range $0.05,1.3]~\mathrm{\AA }^{-1}$.
At $T=0.5$ K and 10 K, we observe quasi-elastic scattering extending up to $\hbar\omega\sim 0.6$ meV, indicative of persistent spin fluctuations above $T_{f}$. To characterize the energy dependence of this quasi-elastic scattering, we adopt a generalized Lorentzian model [45],
$\begin{split}I(\hbar\omega)&=a_{0}\exp\left(-\frac{(\hbar\omega)^{2}}{2\sigma^{2}}\right)+\frac{a_{1}\hbar\omega}{1-\exp(-\hbar\omega/k_{B}T)}\\ &\quad\times\left(\frac{\Gamma}{(\hbar\omega)^{2}+\Gamma^{2}}\right)^{\epsilon(\omega)},\end{split}$ (2)
where the first Gaussian term accounts for elastic contributions including nuclear scattering, static magnetic correlations, and residual background (see Fig. 5(i,ii,iii)). The Gaussian width $\sigma\approx 0.0087$ meV is fixed by the instrumental resolution. The second term describes the quasi-elastic magnetic scattering. The phenomenological exponent $\epsilon(\omega)$ is introduced to capture the high-energy tails, defined as $\epsilon(\omega)=1$ for $|\omega|\leq\Gamma$ and $\epsilon(\omega)=1+\alpha(|\omega|-\Gamma)/\Gamma$ for $|\omega|>\Gamma$ [45]. The best-fit results are shown as dashed lines in Fig. 5(b,c), yielding characteristic energy widths of $\Gamma=0.062(3)$ meV at $T=0.5$ K and $\Gamma=0.10(1)$ meV at $T=10$ K. These correspond to timescales $\tau_{\mathrm{fast}}\sim h/\Gamma$ of approximately 0.07 ns and 0.04 ns, respectively. We interpret these timescales as signatures of the “fast” spin dynamics predicted for QSI systems [36]. In both fits, small values of the exponent $\alpha$ in $\epsilon(\omega)$ were obtained: $\alpha=0.13(2)$ at $T=0.5$ K and $\alpha=0.05(13)$ at $T=10$ K.
At $T=0.05$ K, below $T_{f}$, the magnetic excitation spectrum undergoes a pronounced transformation: quasi-elastic spectral weight in the $[0,0.2]$ meV range is significantly suppressed, and a gapped excitation emerges. This excitation is modeled using the form
$I(\hbar\omega)$ $=a_{0}\exp\left(-\frac{(\hbar\omega)^{2}}{2\sigma^{2}}\right)+\frac{a_{1}\Gamma}{(\hbar\omega-E_{\mathrm{gap}})^{2}+\Gamma^{2}},$ (3)
centered at $E_{\mathrm{gap}}=0.23(1)$ meV ($\sim 2.67$ K) with a linewidth of $\Gamma=0.11(2)$ meV, , in good agreement with the activation energy $E_{a}\approx 2.7$ K from a.c. susceptibility. By integrating the $a_{1}$ components of Eq. 2 and 3, we find that the integrated inelastic intensity at $T=0.05$ K and the quasi-elastic intensity at $T=0.5$ K account for approximately 40% and 80%, respectively, of the total quasi-elastic spectral weight observed at $T=10$ K. To satisfy the total-moment sum rule, we expect that the remaining spectral weight lies underneath the elastic line and corresponds to the slow timescale $\tau_{\mathrm{slow}}$ accessible only to a.c. susceptibility measurements.
We now consider the wavevector dependence of the inelastic scattering. Constant-energy slices collected at $T=0.05$ K and 0.5 K for various energy transfers $\hbar\omega$ are shown in Fig. 5(d–j), using incident energy $E_{i}=3.32$ meV. For each panel, the corresponding $T=10$ K measurement has been subtracted to eliminate non-magnetic contributions. For Fig. 5(f–j), the energy windows were chosen to minimize elastic contributions within the instrumental resolution. Additional details on background subtraction [Fig. A8] and resolution effects [Fig. A11] are provided in Appendix F.
At $T=0.5$ K, we observe a characteristic “pinch-point” pattern in the elastic channel, integrated over $\hbar\omega\in[-0.2,0.2]$ meV [Fig. 5(d)]. Upon cooling into the spin-frozen state at $T=0.05$ K, the $\bm{Q}$-dependence of the elastic scattering changes dramatically, as shown in Fig. 5(e). The resulting pattern is reminiscent of the magnetic diffuse scattering observed in CSI materials with strong dipolar interactions such as $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$[46, 47] and $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$[48, 49], but deviates even more from the ideal pinch-point structure. Instead, the scattering becomes concentrated near reciprocal space positions associated with a $\bm{q}=(100)$ propagation vector, suggesting a tendency towards incipient $\bm{q}=(100)$ order as the system undergoes freezing. The $\bm{Q}$-dependence of the gapped excitation observed at $T=0.05$ K [Fig. 5(a)] is revealed in Fig. 5(g), showing a starfish-shaped distribution reminiscent of a smeared pinch-point pattern, similar to inelastic magnetic features previously observed in several QSI candidates [23, 50, 22, 27]. A similar $\bm{Q}$-dependence is observed for the quasielastic scattering at $T=0.5$ K over the same energy range $\hbar\omega\in[0.2,0.6]$ meV [Fig. 5(f)], although the features appear more diffuse. For higher energy transfers $\hbar\omega\geq 0.6$ meV [Fig. 5(h, j)], no significant $\bm{Q}$-dependence is observed. Additional data are presented in Figs. A7,  A9, and A10.
IV Discussion
Our $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples exhibit magnetic properties consistent with a QSI state at finite temperature $T=0.5$ K. However, the observation of a full spin-freezing transition at $T_{f}=0.15$ K indicates that our $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  sample deviates from ideal QSI behavior and enters a different phase as $T\rightarrow$ 0. This deviation is presumably driven by disorder, which dominates the low-temperature magnetic properties.
A key open question concerns the microscopic nature of the spin-frozen state and the evolution of the neutron spectra as $T_{f}$ is crossed. We propose two mechanisms that may account for this evolution. The first mechanism involves the temperature dependence of CEF transitions between split singlets at sites with moderate transverse field $h_{i}\sim J_{1}$; a mean-field model of this effect is given in Appendix A. As the temperature drops below $T_{f}$, each pseudospin experiences an increasing exchange field from neighboring spins. This enhanced exchange field favors alignment of the pseudospin along the local $\hat{\bm{z}}$ direction. As a result, the spectral weight associated with sites where $h_{i}\sim J_{1}$ is transferred from inelastic CEF transitions to the elastic line, accompanied by an upward shift in the lower bound of the CEF excitation energies [Fig. A1]. However, this mechanism alone is unlikely to account for the full $\sim 40\%$ transfer of magnetic spectral weight that we observed, since once the spin-ice state is established, further cooling should not greatly enhance the exchange field.
The second mechanism assumes that the spin-freezing transition is driven by the spatial variance of the transverse field, $\delta h$, exceeding a critical threshold $\delta h_{c}$ [29, 30]. In this regime, certain ring-exchange terms become significantly enhanced relative to others, favoring the condensation of vison excitations. We interpret the observed suppression of quasielastic scattering below $T_{f}$ as a signature of vison condensation, which suppresses the low-energy quantum fluctuations characteristic of a quantum spin ice (QSI). Theoretically, vison condensation has also been predicted to induce $(100)$ magnetic order [6], which indeed appears incipient in the frozen state [Fig. 5(e)].
[…]
Our results mark an important step toward realizing a U(1) QSL state, as they indicate that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ lies in close proximity to the ideal U(1) QSL in two key respects. First, in the temperature range above $T_{f}$ but below the onset of spin-ice correlations at $\sim$1 K, $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ exhibits hallmarks of a QSL state: the characteristic wavevector dependence of the diffuse scattering [40] and the presence of two well-separated magnetic timescales [36]. Second, the spin-frozen state that develops below $T_{f}$ appears to emerge from the QSL via a confinement transition, in contrast to floating-zone–grown $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ samples [21, 23, 22], which remain in a paramagnetic regime separated from the QSI phase by a first-order transition.
V Conclusion
[…]
Acknowledgements
We thank O. Benton for insightful discussions. B.R.O. (synthesis, bulk property measurements, and analysis) gratefully acknowledges support from the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. A portion of this research was conducted using resources at the Spallation Neutron Source and the High Flux Isotope Reactor, DOE Office of Science User Facilities operated by Oak Ridge National Laboratory, under Proposals IPTS-32079, 31234, and 34042. S.D.W. acknowledges support from the DOE Office of Basic Energy Sciences, Division of Materials Sciences and Engineering, under Grant No. DE-SC0017752. Part of this work utilized facilities at the UC Santa Barbara NSF Quantum Foundry, supported through the Q-AMASE-i program under award DMR-1906325. We also thank the SNS X-ray Laboratory at ORNL for access to the Laue instrument used in sample coalignment. We are grateful to Matthew Powell and Joseph W. Kolis (Clemson University) for providing the La2Sn2O7 single crystal, which was used as a non-magnetic background in the heat capacity measurements reported in this work.
This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the U.S. Department of Energy (DOE). The U.S. government retains, and the publisher—by accepting the article for publication—acknowledges that the U.S. government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or to allow others to do so, for U.S. government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-publicaccess-plan).
Appendix A Model for the gapped excitation in spin-frozen state
Figure A1: (a) Evolution of the neutron scattering cross-sections for the elastic channel ($I_{\mathrm{el}}$) and the inelastic channel ($I_{\mathrm{inel}}$), centered at energy transfer $\Delta_{\mathrm{gap}}$, as a function of the ratio between the (uniform) transverse field $h_{i}$ and the mean-field spin-flip energy $E_{J}$ in the classical spin-ice limit, calculated using Eq. 6. (b) Simulated elastic and inelastic magnetic scattering intensities based on the mean-field model in Eq. 4, assuming a half-Gaussian distribution of transverse fields $h_{i}$ as defined in Eq. 7 with $\sigma=0.14$ meV, for two representative values of $E_{J}$. For $E_{J}=0.13$ meV (0.065 meV), the inelastic contribution $I_{\mathrm{inel}}$ constitutes approximately 34% (40%) of the total magnetic scattering intensity, $I_{\mathrm{el}}+I_{\mathrm{inel}}$. The elastic line is convolved with a full width at half maximum (FWHM) of 0.02 meV, consistent with the instrumental resolution of the $E_{i}=0.99$ meV neutron spectroscopy setup. The peak of the elastic intensity is shown in inset (i).
As discussed in Section IV, the spin-frozen phase below $T_{f}\approx 0.15$ K hosts a gapped excitation at $\hbar\omega=0.23$ meV with a starfish-like $\bm{Q}$ dependence, reminiscent of inelastic features reported in other QSI candidates [23, 50, 22, 27]. In $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$, analogous excitations were attributed to CEF transitions between singlets of a split non-Kramers doublet, with the $\bm{Q}$ dependence shaped by finite nearest-neighbor correlations $J_{1}$ [22, 30]. Within a mean-field framework [23, 24], the elastic and inelastic spectral weights scale as $|\cos\theta_{i}|^{2}$ and $|\sin\theta_{i}|^{2}$, respectively, where $\theta_{i}$ is the angle between the pseudospin (defined in Eq. 1) and the local $\hat{\bm{z}}$ axis. Our finding that the inelastic channel carries only $\sim$40% of the total weight indicates a moderate deviation of the pseudospins from the Ising axis and suggests that the system retains features of the paramagnetic phase in the phase diagram [Fig. 1].
To characterize further the behavior of these CEF excitations, we follow Ref. [24] and introduce a mean-field model in which disorder in Ising pyrochlores with a non-Kramers ground-state doublet is represented by a site-dependent transverse field $h_{i}$ acting on each Pr3+ ion [29, 30]. The system is described by the Hamiltonian given by Eq. 4.
When $h_{i}$ is finite, the pseudo-spin $\sigma_{i}$ at site $i$ is no longer aligned strictly along the local $\langle 111\rangle$ $\hat{\bm{z}}$ direction. Following Refs. [23, 24], we consider a local mean-field Hamiltonian for site $i$,
$\mathcal{H}_{\mathrm{mf}}=\begin{pmatrix}0&-h_{i}\\ -h_{i}&E_{J}\end{pmatrix}$, (4)
[…]
where $a_{1}$, $b_{1}$, $a_{2}$, and $b_{2}$ are normalized coefficients, satisfying $|a_{1}|^{2}+|b_{1}|^{2}=1$ and $|a_{2}|^{2}+|b_{2}|^{2}=1$. Compared to the convention adopted in Ref. [23], where a local spherical angle $\theta_{i}$ is defined between the pseudospin and the $\hat{\bm{z}}$ direction, the parameters used here are related by: $a_{1}\simeq\sin(\theta_{i}/2)$, $b_{1}\simeq\cos(\theta_{i}/2)$, $a_{2}\simeq\cos(\theta_{i}/2)$, and $b_{2}\simeq-\sin(\theta_{i}/2)$.
Since neutrons primarily couple to the $\hat{\bm{z}}-$component of the real spin for systems with non-Kramers ground state doublet [7], and the projected spin operator $J_{z}$ within the non-Kramers doublet is proportional to the pseudo-spin operator $\sigma_{z}$ [23], the neutron scattering intensities in the elastic and inelastic channels can be expressed as:
[…]
In the classical spin ice limit ($h_{i}=0$), we have $a_{2}=b_{1}=1$, meaning the ground state is fully aligned along $\hat{\bm{z}}$ and follows the “two-in, two-out” ice rule per tetrahedron. In this case, $I_{\mathrm{inel}}=0$, and all magnetic spectral weight resides at the elastic line. In the opposite extreme ($h_{i}\gg J_{1}$), the pseudo-spin lies nearly perpendicular to the local easy axis; $a_{1}\approx b_{2}\approx 1$, leading to a singlet ground state with vanishing elastic scattering ($I_{\mathrm{el}}\rightarrow 0$) and all spectral weight transferred to the inelastic channel at energy $\Delta_{\mathrm{gap}}$.
The evolution of $I_{\mathrm{el}}$ and $I_{\mathrm{inel}}$ with fixed $E_{j}$ and increasing $h_{i}$ is shown in Fig. A1(a). To model the distribution of transverse fields, we assume a half-Gaussian distribution ($p(h)=0$ for $h<0$),
$p(h)=\frac{2}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{h^{2}}{2\sigma^{2}}\right),\quad h\geq 0,$ (7)
with $\sigma=0.14$ meV. The resulting simulated inelastic spectrum is shown in Fig. A1(b) for two selected values of $E_{J}$, illustrating that approximately 34% (40%) of the total magnetic scattering intensity, $I_{\mathrm{el}}+I_{\mathrm{inel}}$, originates from the inelastic component for $E_{J}=0.13$ meV ($E_{J}=0.065$ meV). The choice of $E_{J}\approx 0.13$ meV corresponds to the fitted nearest-neighbor interaction $J_{1}=0.38(1)$ K, based on magnetic diffuse scattering data. The elastic contribution is convolved with the instrumental energy resolution (FWHM = 0.02 meV), consistent with the $E_{i}=0.99$ meV setup used in our neutron spectroscopy experiment.As discussed previously for the constant $h_{i}$ case, increasing $E_{J}$ leads to an inelastic spectrum characterized by a larger gap $\Delta_{\mathrm{gap}}$ and reduced spectral weight. The simulated inelastic response exhibits a sharp onset at $E_{J}$, followed by a long tail extending to higher energies. This behavior is expected, as the inelastic matrix element $I_{\mathrm{inel}}$ increases with the ratio $h_{i}/E_{J}$. Importantly, the high-energy tail is relatively insensitive to variations in $E_{J}$, as it arises from sites where the transverse field dominates, i.e., $h_{i}\gg E_{J}$. In practice, the energy dependence of $I_{\mathrm{inel}}$ is further influenced by the specific shape of the transverse field distribution $p(h)$ and by the dispersion of gapped excitations, which lies beyond the scope of the mean-field approximation.
Figure A2: (a) Photograph of a portion of the co-aligned single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  array (total mass $\sim 1.27$ g), with the vertical axis approximately aligned along the $1\bar{1}0]$ crystallographic direction. (b,c) Photographs of the crushed single-crystal sample used for the neutron powder diffraction experiment on the Nanoscale-Ordered Materials Diffractometer (NOMAD), exhibiting a dark red color. (d,e) Optical microscopy images zooming in on a representative crystal affixed to the copper plate within the co-aligned array shown in panel (a); the scale bar is shown in panel (e). (f) Neutron diffraction pattern of the co-aligned array collected at $T=0.3$ K with an incident wavelength of $\lambda=1.5$ Å on the wide-angle single-crystal diffractometer WAND2. The data indicate a single, slightly broadened crystal grain with an estimated mosaicity of $\sim 5^{\circ}$ (FWHM), as shown in (g–j). (g–j) Transverse cuts at the Bragg peaks $\bm{Q}=(880)$ and $\bm{Q}=(008)$ along the in-plane directions ($[001]$ and $[110]$, respectively) and the out-of-plane direction $K\bar{K}0]$ at $T=0.3$ K. The horizontal axes are plotted in angular units $\psi_{\mathrm{in}}$ (in-plane), $\phi_{\mathrm{out}}$ (out-of-plane) by normalizing to $|\bm{Q}|$. Data are integrated within $\pm 0.08~\mathrm{\AA }^{-1}$ in the perpendicular $\bm{Q}$-directions.
Figure A3: (a–j) Frequency dependence of the real part of the ac susceptibility $\chi^{\prime}$ for various temperatures in the absence of an external dc field , fits to the damped harmonic oscillator model in Eq. 8.
Appendix B Subtraction of nuclear Schottky anomaly in heat capacity
We analyzed the nuclear Schottky anomaly in $C_{p}$ using two models: (1) the analytical model of Kimura et al. [21], with $\mu_{\mathrm{hyp}}^{\mathrm{Pr}}=1.9~\mathrm{\upmu_{B}}$ and no scaling factor, as in Ref. [21, 35]; and (2) a statistical model (Gopal model) [43] treating the system as a multilevel scheme with evenly spaced energy levels, using a level spacing of $\epsilon_{r}=0.011$ meV. Both models successfully reproduce the broad maximum in $C_{p}$ near 0.1 K. We note that our heat capacity data and the parameters obtained from model (1) are very similar to those of $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ sample 1 reported in Ref. [44]. The temperature dependence of the magnetic entropy $S_{m}$, shown in Fig. 4(ii), was determined by subtracting both the Gopal-model nuclear contribution and the lattice contribution (estimated from nonmagnetic La2Sn2O7) from the measured $C_{p}$, and integrating the resulting magnetic contribution $C_{m}/T$. We emphasize that subtraction using either the Kimura or the Gopal model yields essentially identical results for the magnetic entropy.
Appendix C Additional information for $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ samples
[…]
Magnetization versus field measurements at $T=2$ K on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples with masses of approximately 5 mg are shown in Fig. A4(c). The induced magnetic moments $M$ at the maximum applied field $\mu_{0}H=7$ T are measured to be $1.64~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, $1.33~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, and $1.19~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$ for fields applied along the $[100]$, $[111]$, and $[110]$ directions, respectively. For comparison, the expected saturated moments in a pyrochlore lattice with Ising anisotropy along the local $\langle 111\rangle$ directions and effective ferromagnetic (FM) coupling between nearest neighbors are $g_{J}J/\sqrt{3}$ for the $[100]$ direction, $g_{J}J/2$ for $[111]$, and $g_{J}J/\sqrt{6}$ for $[110]$ [51]. Both the observed ordering of the induced moments (from largest to smallest) and their relative ratios are consistent with the predictions for a magnetic pyrochlore with easy-axis FM anisotropy. Quantitative deviations may arise from slight misalignments during sample mounting, mass uncertainty due to residual impurity phases, or the fact that full saturation may not have been achieved at $\mu_{0}H=7$ T. The temperature dependence of the dc magnetic susceptibility was measured with the magnetic field applied along the $[100]$, $[111]$, and $[110]$ crystallographic directions using the same single-crystal samples. To obtain a representative bulk behavior, we compute a powder-averaged dc susceptibility for a cubic system using the expression $\chi_{\mathrm{powder}}^{\mathrm{DC}}=(3\chi_{100}^{\mathrm{DC}}+6\chi_{110}^{\mathrm{DC}}+4\chi_{111}^{\mathrm{DC}})/13$ [52], and plot the results in Fig. A4(d,e). This powder-averaged dc susceptibility is also used to normalize the real part of the ac susceptibility $\chi^{\prime}$ presented in Fig. 5(a). A Curie–Weiss fit to the high-temperature regime yields a Curie–Weiss temperature of $\theta_{\mathrm{CW}}=-0.23(2)$ K and an effective moment of $\mu_{\mathrm{eff}}=2.78(1)~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, in good agreement with the powder measurements reported in Ref. [35].
To extract the “slow” timescale of spin dynamics [36] from the real part $\chi^{\prime}$ of the ac susceptibility, we employ a damped harmonic oscillator model [42, 35]:
$\chi^{\prime}(f)=\chi_{S}+\frac{\chi_{T}-\chi_{S}}{1+(2\pi f\tau)^{2}},$ (8)
where $f$ is the driving frequency, $\tau$ is the characteristic spin correlation time, $\chi_{S}$ is the adiabatic (high-frequency) susceptibility, and $\chi_{T}$ is the isothermal (low-frequency) susceptibility. The details of the fits at each temperature are shown in Fig. A3(a–j) and the results are presented in Fig. 4(i).
Appendix D Modeling of magnetic diffuse-scattering data
Figure A4: (a) Goodness-of-fit map ($\chi^{2}$) as a function of the ring-exchange parameter $g$ (defined in Eq. 11) and the theoretical tuning parameter $\mu/g$ for the quantum spin ice model. The quantum limit corresponds to $\mu/g=0$, where the best-fit value is $g=0.075(1)$ K. (b) Goodness-of-fit map ($\chi^{2}$) versus the nearest- and next-nearest-neighbor exchange interactions $J_{1}^{\mathrm{ex}}$ and $J_{2}^{\mathrm{ex}}$ (defined in Eq. 13). The best-fit parameters are $J_{1}^{\mathrm{ex}}=0.25(1)$ K and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K. (c) Magnetization versus field measured at $T=2$ K on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples with masses of approximately 5.0 mg, 6.3 mg, and 5.6 mg for magnetic fields applied along the $[100]$, $[111]$, and $[110]$ directions, respectively. (d,e) Temperature dependence of powder-averaged dc susceptibility measured on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, shown alongside SPINTERACT [41] simulations using the best-fit parameters from panel (b).
Figure A5: Illustration of the sign convention for the fitted interaction parameters from the classical model using the SPINTERACT program [41]. An antiferromagnetic nearest-neighbor interaction $J_{1}>0$ in the local frame (defined in Eq. 9) favors the canonical “two-in–two-out” spin configuration (red arrows) on each tetrahedron. A ferromagnetic next-nearest-neighbor interaction $J_{2}<0$ in the local frame favors a “loop” structure, where spins on three next-nearest-neighbor sites around a hexagonal plaquette align either all clockwise or all counterclockwise relative to the plaquette center.
To model the magnetic diffuse scattering observed in single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, we use the SPINTERACT software suite [41]. For quantum and classical models, we write the Hamiltonian in terms of unit Pauli matrices or unit-length Ising variables, respectively, corresponding to an effective spin-$1/2$ on each Pr3+ site with an effective $g$-factor $g_{\parallel}=5.17$ along the local $\langle 111\rangle$ direction [28]. The local coordinate frames $(\hat{\mathbf{a}}_{i},\hat{\mathbf{b}}_{i},\hat{\mathbf{e}}_{i})$ are defined as in Ref. [8]:
$\begin{split}\hat{\mathbf{e}}_{0}&=\frac{1}{\sqrt{3}}(1,1,1),~~~~~~~\hat{\mathbf{a}}_{0}=\frac{1}{\sqrt{6}}(-2,1,1),\\ \hat{\mathbf{e}}_{1}&=\frac{1}{\sqrt{3}}(1,-1,-1),\quad\hat{\mathbf{a}}_{1}=\frac{1}{\sqrt{6}}(-2,-1,-1),\\ \hat{\mathbf{e}}_{2}&=\frac{1}{\sqrt{3}}(-1,1,-1),\quad\hat{\mathbf{a}}_{2}=\frac{1}{\sqrt{6}}(2,1,-1),\\ \hat{\mathbf{e}}_{3}&=\frac{1}{\sqrt{3}}(-1,-1,1),\quad\hat{\mathbf{a}}_{3}=\frac{1}{\sqrt{6}}(2,-1,1),\end{split}$ (9)
corresponding to the four Pr3+ ions located at $\hat{\mathbf{r}}_{0}=(0.5,0.5,0.5)$, $\hat{\mathbf{r}}_{1}=(0.5,0.75,0.75)$, $\hat{\mathbf{r}}_{2}=(0.75,0.5,0.75)$, and $\hat{\mathbf{r}}_{3}=(0.75,0.75,0.5)$ in the face-centered cubic unit cell. A magnetic background following the form factor $|F(Q)|^{2}$ for Pr3+ is subtracted to simulate the experimental high-temperature background removal. The SPINTERACT program minimizes the goodness-of-fit metric $\chi^{2}$, defined as
$\begin{split}\chi^{2}&=\sum_{d}W_{d}\sum_{i\in d}\left(\frac{I_{i}^{\text{data}}-s_{d}I_{i}^{\text{calc}}}{\sigma_{i}}\right)^{2},\\ s_{d}&=\frac{\sum_{i}I_{i}^{\text{data}}I_{i}^{\text{calc}}/\sigma_{i}^{2}}{\sum_{i}(I_{i}^{\text{calc}})^{2}/\sigma_{i}^{2}},\end{split}$ (10)
where $d$ indexes a particular dataset with weight $W_{d}$, $I_{i}^{\text{data}}$ and $I_{i}^{\text{calc}}$ denote the measured and calculated intensities for the $i$-th data point, respectively, and $\sigma_{i}$ is the corresponding experimental uncertainty. The scale factor $s_{d}$ for each dataset is determined via the linear least-squares expression given above. For the quantum model, we adopt the theoretical results of Ref. [40], where the goodness-of-fit $\chi^{2}$ is plotted against the ring-exchange parameter $g$ in the effective Hamiltonian:
$\mathcal{H}_{\text{ring}}=-\sum_{\hexagon}g\left(\sigma^{+}_{i}\sigma^{-}_{j}\sigma^{+}_{k}\sigma^{-}_{l}\sigma^{+}_{m}\sigma^{-}_{n}+\text{H.c.}\right).$ (11)
Here, $\mu$ is a theoretical tuning parameter, with $\mu/g=0$ corresponding to the quantum spin ice (QSI) limit and $\mu/g=1$ to the classical limit. This model assumes perturbative interactions with nearest-neighbor exchange $J_{1}\rightarrow\infty$. Best-fit analysis yields $g=0.075(1)$ K in the QSI limit ($\mu/g=0$). In the classical limit ($\mu/g=1$), the scattering profile is independent of $g$, and the result corresponds to the $J_{1}$-only classical spin ice scenario shown in Fig. 3(c). The decreasing trend of $\chi^{2}$ with increasing $g$ in this regime is attributed to numerical artifacts.
For the classical model, we employ Onsager reaction-field theory, incorporating long-range dipolar interactions $D_{\mathrm{dip}}\approx 0.012$ K, following the convention of Ref. [41]. The total interaction $J_{n}$ between $n$-th neighbor $\sigma_{i}^{z}$ terms in Eq. 1 thus includes both dipolar and exchange contributions,
$J_{n}=D_{n}+J_{n}^{\mathrm{ex}}.$ (12)
Further-neighbor interactions, in particular $J_{2}$ and $J_{3}$, have also been invoked to explain the magnetic correlations observed below the spin-freezing transition in $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$ [48, 49]. The Hamiltonian, excluding disorder terms, is then:
$\begin{split}\mathcal{H}&=\sum_{\langle ij\rangle_{1}}(J_{1}^{\mathrm{ex}}+D_{1})\sigma_{i}^{z}\sigma_{j}^{z}+\sum_{\langle ij\rangle_{2}}(J_{2}^{\mathrm{ex}}+D_{2})\sigma_{i}^{z}\sigma_{j}^{z}\\ &+\text{Further-neighbor terms}\end{split}$ (13)
where $D_{1}=0.13$ K and $D_{2}=-0.03$ K are the effective dipolar couplings for nearest and next-nearest neighbors, respectively. All parameters are expressed in the local coordinate frames defined in Eq. 9 and the long-ranged nature of the dipolar interaction was implemented using Ewald summation. The magnetic diffuse scattering pattern alone does not effectively constrain the magnitude of $J_{1}^{\mathrm{ex}}$, so we complement the fitting with powder-averaged d.c. susceptibility data ($W_{d}=1$ for all data sets). A $\chi^{2}$ map of the fits to d.c. susceptibility as a function of $J_{1}^{\mathrm{ex}}$ and $J_{2}^{\mathrm{ex}}$ is shown in Fig. A4(b), with the optimal parameters found to be $J_{1}^{\mathrm{ex}}=0.25(1)$ K and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K. The sign convention for the fitted interaction parameters is illustrated in Fig. A5. An antiferromagnetic nearest-neighbor interaction $J_{1}=J_{1}^{\mathrm{ex}}+D_{1}>0$ between pseudo-spins in the local frame favors the canonical “two-in, two-out” spin configuration on each tetrahedron. In contrast, a ferromagnetic next-nearest-neighbor interaction $J_{2}=J_{2}^{\mathrm{ex}}+D_{2}<0$ in the local frame favors a “loop” structure, where spins on three next-nearest-neighbor sites around a hexagonal plaquette align either all clockwise or all counterclockwise relative to the plaquette center. The corresponding fits to the susceptibility data are shown in Fig. A4(d,e).
Appendix E Powder neutron diffraction refinements
Rietveld refinement results are provided in Table 1c(a,b,c), and are shown in Fig. A6, assuming a stoichiometric structure, since refinements of the expected site-disorder models discussed in the main text yielded $\lesssim 1$% deviations from ideal site occupancies.
Figure A6: Rietveld refinement of neutron powder diffraction patterns measured on the Nanoscale-Ordered Materials Diffractometer (NOMAD) at 100 K (a) and 300 K (b). Observed data are shown as black points, calculated patterns as red lines, Bragg reflection positions as green tick marks, and the difference between observed and calculated intensities as blue lines.
Table 1: Atomic positions and anisotropic displacement parameters of crushed single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  collected on the Nanoscale-Ordered Materials Diffractometer (NOMAD), assuming full stoichiometry.
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.00070(7) 0.00070(7) 0.00070(7) 0.00004(7) 0.00004(7) 0.00004(7)
Pr (16d) 0.5 0.5 0.5 0.00257(9) 0.00257(9) 0.00257(9) -0.00047(11) -0.00047(11) -0.00047(11)
O (8b) 0.375 0.375 0.375 0.00263(10) 0.00263(10) 0.00263(10) 0 0 0
O (48f) 0.331480(18) 0.125 0.125 0.00266(10) 0.00334(8) 0.00334(8) 0 0 0.00139(9)
(a) Rietveld refinement from data collected at $T=100$ K
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.00180(9) 0.00180(9) 0.00180(9) 0.00028(8) 0.00028(8) 0.00028(8)
Pr (16d) 0.5 0.5 0.5 0.00507(13) 0.00507(13) 0.00507(13) -0.00102(12) -0.00102(12) -0.00102(12)
O (8b) 0.375 0.375 0.375 0.00400(13) 0.00400(13) 0.00400(13) 0 0 0
O (48f) 0.331219(19) 0.125 0.125 0.00452(12) 0.00537(10) 0.00537(10) 0 0 0.00252(10)
(b) Rietveld refinement results from data collected at $T=300$ K
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.00167(4) 0.00167(4) 0.00167(4) -0.00022(7) -0.00022(7) -0.00022(7)
Pr (16d) 0.5 0.5 0.5 0.00285(7) 0.00285(7) 0.00285(7) -0.00027(1) -0.00027(1) -0.00027(1)
O (8b) 0.375 0.375 0.375 0.00353(10) 0.00353(10) 0.00353(10) 0 0 0
O (48f) 0.331348(52) 0.125 0.125 0.00465(12) 0.00396(8) 0.00396(8) 0 0 0.00144 (9)
(c) PDF refinement from data collected at $T=100$ K
Table 2: Atomic positions and anisotropic displacement parameters of single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  at various temperatures and measurement methods, with occupancy of Pr site indicated in the captions.
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.00096(7) 0.00096(7) 0.00096(7) -0.00004(7) -0.00004(7) -0.00004(7)
Pr (16d) 0.5 0.5 0.5 0.00179(9) 0.00179(9) 0.00179(9) -0.00031(10) -0.00031(10) -0.00031(10)
O (8b) 0.375 0.375 0.375 0.00278(10) 0.00278(10) 0.00278(10) 0 0 0
O (48f) 0.331509(17) 0.125 0.125 0.00308(10) 0.00351(8) 0.00351(8) 0 0 0.00137(9)
(a) Refinement results from data collected at $T=100$ K on NOMAD with Pr-site occupancy $95.30(16)\%$.
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.00214(9) 0.00214(9) 0.00214(9) 0.00016(8) 0.00016(8) 0.00016(8)
Pr (16d) 0.5 0.5 0.5 0.00408(12) 0.00408(12) 0.00408(12) -0.00079(12) -0.00079(12) -0.00079(12)
O (8b) 0.375 0.375 0.375 0.00422(13) 0.00422(13) 0.00422(13) 0 0 0
O (48f) 0.331219(19) 0.125 0.125 0.00503(12) 0.00560(10) 0.00560(10) 0 0 0.00248(10)
(b) Refinement results from data collected at $T=300$ K on NOMAD with Pr-site occupancy $95.30(16)\%$.
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.0031(5) 0.0031(5) 0.0031(5) 0.00005(4) 0.00005(4) 0.00005(4)
Pr (16d) 0.5 0.5 0.5 0.0044(4) 0.0044(4) 0.0044(4) -0.00063(4) -0.00063(4) -0.00063(4)
O (8b) 0.375 0.375 0.375 0.0071(8) 0.0071(8) 0.0071(8) 0 0 0
O (48f) 0.33155(16) 0.125 0.125 0.0064(10) 0.0075(6) 0.0075(6) 0 0 0.0020(7)
(c) Refinement results from data collected at $T=300$ K with X-rays, Pr-site occupancy $98.84(14)\%$.
[…]
Atom (Wyckoff) $x$ $y$ $z$ $u_{11}$ $u_{22}$ $u_{33}$ $u_{12}$ $u_{13}$ $u_{23}$
Sn (16c) 0 0 0 0.0022(6) 0.0022(6) 0.0022(6) 0.00006(6) 0.00006(6) 0.00006(6)
Pr (16d) 0.5 0.5 0.5 0.0039(5) 0.0039(5) 0.0039(5) -0.00065(5) -0.00065(5) -0.00065(5)
O (8b) 0.375 0.375 0.375 0.0057(9) 0.0057(9) 0.0057(9) 0.000 0.000 0.000
O (48f) 0.3312(2) 0.125 0.125 0.0032(11) 0.0061(7) 0.0061(7) 0.000 0.000 0.0013(8)
(d) Refinement results from data collected at $T=300$ K with X-rays, Pr-site occupancy $100\%$.
Figure A7: (a,b) Powder-averaged intensity as a function of energy transfer $\hbar\omega$ at $T=0.05$, 0.5, and 10 K, measured with incident energies $E_{i}=0.99$ meV and 3.32 meV, respectively. Data are integrated along $K\bar{K}0]$ with $K\in[-0.1,0.1]$ rlu ([0.15,0.15][-0.15,0.15] rlu) and over $Q\in[0.05,1.3]~\mathrm{\AA }^{-1}$ ($Q\in[0.15,2.5]~\mathrm{\AA }^{-1}$ ) within the horizontal $(HHL)$ plane for $E_{i}=0.99$ meV ($E_{i}=3.32$ meV). Insets (i,ii) show the corresponding elastic intensities. (c-h) False-color maps of powder-averaged intensity as functions of momentum transfer $|Q|$ and energy transfer $\hbar\omega$. Panels (c-e) correspond to $E_{i}=0.99$ meV, while panels (f-h) correspond to $E_{i}=3.32$ meV.
Figure A8: (a) Determination of the shoulder-like background feature at $\hbar\omega\sim 0.1$ meV observed in Fig. A7(a), measured with $E_{i}=0.99$ meV. The black curve shows the raw inelastic scattering data on the energy-loss side ($\hbar\omega>0$) at $T=10$ K. The red curve represents the calculated energy-loss intensity obtained from the energy-gain side via the detailed balance relation $I(|\hbar\omega|)=I(-|\hbar\omega|)\exp(|\hbar\omega|/k_{B}T)$. The green curve shows the difference between the measured (black) and calculated (red) curves, representing the estimated temperature-independent background. (b,i) Same data as Fig. A7(a), but with the background determined in (a) subtracted for all three temperatures.
Figure A9: (a–c) False-color maps of the scattering intensity as a function of momentum transfer along the $[00L]$ direction and energy transfer $\hbar\omega$, collected at $T=0.05$ K, $T=0.5$ K, and $T=10$ K, respectively. (d–f) Corresponding false-color maps of the scattering intensity along the $[HHH]$ direction as a function of energy transfer $\hbar\omega$, at the same temperatures. All data were collected with incident energy $E_{i}=0.99$ meV and integrated over $\pm 0.15$ rlu in the directions perpendicular to the respective cut directions.
Figure A10: (a–h) Constant-energy $Q$-maps measured at $T=0.5$ K (a,c,e,f) and $T=0.05$ K (b,d,f,h) using incident energy $E_{i}=0.99$ meV (elastic resolution FWHM $\approx 0.02$ meV), following subtraction of the $T=10$ K background. Data are integrated along $K\bar{K}0]$ with $K\in[-0.1,0.1]$ rlu, and over the energy ranges $\hbar\omega\in[-0.05,0.05]$ meV (a, b), $[0.05,0.15]$ meV (c, d), $[0.15,0.25]$ meV (e, f), and $[0.25,0.35]$ meV (g, h). No symmetrization is applied. Color scales are centered on the mean intensity of each panel. Energy-gain and energy-loss contributions are combined in (b-d) panels. (j) Powder-averaged intensity as a function of energy transfer $\hbar\omega$ at $T=0.05$ K for scattering near the horizontal $[HHL]$ plane ($K\leq 0.1$ rlu in the $K\bar{K}0]$ direction), measured with $E_{i}=0.99$ meV. The green solid curve represents intensity in this $[HHL]$ plane integrated within $\pm 0.1$ rlu around the (111) Bragg peak, scaled to match the elastic intensity shown in inset (i). This curve serves as an in-situ measure of the instrumental energy resolution.
When analyzing the data collected on NOMAD, we found that allowing the Pr-site occupancy to vary led to a modest improvement in the refinement, corresponding to an apparent $\sim$4.7% reduction in Pr occupancy. For the 100 K dataset, this yielded $R_{\mathrm{wp}}=3.08\%$, compared with $R_{\mathrm{wp}}=3.15\%$ when full stoichiometry was enforced. The refinement results are summarized in Table 2d(a,b).
[…]
Given the high resolution and good statistics of the single crystal X-ray data, and the presence of significant ($\sim$50%) covariance terms involving the Pr occupancy in the neutron refinements, we believe the actual vacancies (if present) are closer to the 1% level. Fixing the Pr occupancy to unity has a very small effect on the refinement quality ($R_{\mathrm{wp}}$ 2.53% vs 2.59%). Significant levels of Pr vacancies have not been commonly reported in rare-earth pyrochlores, and our results may therefore reflect instrumental or model-dependent effects. Moreover, a large fraction of Pr vacancies would likely necessitate accompanying changes on the oxygen or Sn sites, neither of which was observed in either neutron or X-ray measurements. Given that SCXRD results indicate a nearly fully occupied Pr site, we fix the occupancy to unity and report the corresponding refinement results in Table 2d(d). Further studies will be required to determine whether the apparent deviations in occupancy reflect genuine non-stoichiometry or arise from subtle modeling uncertainties.
Figure A11: (a) Powder-averaged scattering intensity as a function of energy transfer $\hbar\omega$ at $T=0.05$ K, measured near the horizontal $[HHL]$ plane ($K\leq 0.15$ rlu along $K\bar{K}0]$) with incident energy $E_{i}=3.32$ meV. The green solid curve shows the intensity in this $[HHL]$ plane integrated within $\pm 0.15$ rlu around the (111) Bragg peak, scaled to match the elastic peak intensity shown in the inset (i), and serves as an in-situ measure of the instrumental energy resolution. Approximately $99.49\%$ of the integrated intensity of the (111) Bragg peak lies within the energy window [0.2,0.2][-0.2,0.2] meV. (b,c) Constant-energy $\bm{Q}$-maps at $T=0.05$ K, following subtraction of the $T=10$ K background, measured with $E_{i}=3.32$ meV. The data are integrated along $K\bar{K}0]$ with $K\in[-0.15,0.15]$ rlu and over energy ranges $\hbar\omega\in[-0.2,0.2]$ meV (b) and $[0.2,0.6]$ meV (c). No symmetrization is applied. The color scales in panels (b) and (c) are centered on the mean intensity of each panel and span symmetrically about the mean with the same range. In (c), energy-gain and energy-loss contributions are combined. The intensity in (b) is scaled by $1-99.49\%=0.51\%$ to represent the expected elastic leakage into the $[0.2,0.6]$ meV window due to finite resolution, confirming that the observed starfish-shaped $\bm{Q}$-dependence in (c) cannot be attributed to elastic leakage.
Figure A12: Constant-energy $\bm{Q}$-maps at $T=0.5$ K (a,c,e,g,i,k) and $T=0.05$ K (b,d,f,h,j,l), measured with $E_{i}=3.32$ meV (elastic resolution FWHM $\approx 0.1$ meV) after subtraction of the $T=10$ K background. Data are integrated along $K\bar{K}0]$ with $K\in[-0.15,0.15]$ rlu and over the energy ranges marked on the titles of each panel. No symmetrization is applied. Color scales are centered on the mean intensity of each panel.
Appendix F Additional neutron spectroscopy data
In this section, we present additional data and figures from neutron spectroscopy measurements performed on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples. Fig. A7 shows the powder-averaged scattering intensity as a function of energy transfer. For measurements with $E_{i}=0.99$ meV, a gapped inelastic continuum centered at $\hbar\omega\approx 0.2$ meV is observed at $T=0.05$ K, while a broad quasi-elastic continuum extending up to $\hbar\omega\approx 0.6$ meV emerges at $T=0.5$ K. In contrast, the $E_{i}=3.32$ meV data exhibit no discernible additional inelastic features above 0.6 meV. Fig.  A9 presents false-color maps of the scattering intensity as a function of momentum transfer along the $[00L]$ and $[HHH]$ directions and energy transfer, collected with $E_{i}=0.99$ meV. These maps reveal consistent spectral features with those observed in the powder-averaged data shown in Fig. A7(c-e).
F.1 Subtraction of background
In Fig. A7(a), we identify a temperature-independent shoulder-like background feature centered around $\hbar\omega\sim 0.1$ meV, which is absent on the energy-gain side ($\hbar\omega<0$) at $T=10$ K. To remove this feature, we assume the background is temperature independent and there is negligible background at the energy-gain side; this assumption is supported by the observation of negligible signal at $Q>0.2$ Å-1 for $E<0$ at $T=0.05$ K [Fig.A7(c)], at which temperature true inelastic signals will be strongly suppressed for $E<0$. As illustrated in Fig. A8(a), we use the energy-gain side data at $T=10$ K and compute the corresponding energy-loss side ($\hbar\omega>0$) spectrum using the detailed balance relation $I(|\hbar\omega|)=I(-|\hbar\omega|)\exp(|\hbar\omega|/k_{B}T)$ (red line). Here, we neglect the asymmetry in the energy resolution between energy-loss and energy-gain sides, as the energy range is close to the elastic line. The background contribution (green line) is then estimated by subtracting the calculated energy-loss spectrum (red) from the measured data (black). This background is subtracted from the spectra at all three temperatures, yielding the corrected data shown in Fig. A8(b), which is used for the quantitative analysis in Fig. 5(a–c) of the main text.
F.2 Details of constant-energy slices
In Fig. A10(a–h), we present constant-energy $\bm{Q}$-maps similar to those shown in Fig. 5(d–j), but obtained using an incident energy of $E_{i}=0.99$ meV. This configuration offers excellent energy resolution, albeit with limited $Q$-coverage and reduced neutron flux. At $T=0.05$ K, we observe signatures of the incipient “(100)” order, appearing as diffuse scattering centered near $(001)$ and $(110)$ positions [Fig. A10(b)]. However, due to limited statistics and $Q$-range in this configuration, the momentum dependence of the quasi-elastic scattering at $T=0.5$ K and the gapped inelastic features at $T=0.05$ K cannot be resolved clearly.
To overcome this limitation, we employed the $E_{i}=3.32$ meV configuration, which provides a broader $\bm{Q}$-coverage and sufficient flux to resolve the $\bm{Q}$-dependence of the gapped excitation centered at $\hbar\omega=0.23$ meV at $T=0.05$ K. Fig. A11(a) shows the powder-averaged scattering intensity as a function of energy transfer $\hbar\omega$ at $T=0.05$ K, measured near the horizontal $[HHL]$ plane ($K\leq 0.15$ rlu along $K\bar{K}0]$). Also shown is the intensity in this $(HHL)$ plane integrated within $\pm 0.15$ rlu around the (111) Bragg peak, scaled to match the elastic peak shown in Fig. A11(i). The energy profile of the (111) peak serves as an in-situ measure of the instrumental energy resolution, with approximately $99.49\%$ of the total intensity confined within the window $\hbar\omega\in[-0.2,0.2]$ meV, and only $\sim 0.51\%$ lying outside this range. In Fig. A11(b), we present the $\bm{Q}$-map integrated over $\hbar\omega\in[-0.2,0.2]$ meV, scaled by a factor of $\sim 0.51\%$ to estimate the expected elastic leakage into the $[0.2,0.6]$ meV window due to the finite energy resolution. A comparison with the $\bm{Q}$-map integrated over $|\hbar\omega|\in[0.2,0.6]$ meV shown in Fig. A11(c) reveals that the observed starfish-shaped $\bm{Q}$-dependence in this energy range cannot be attributed to elastic leakage. This conclusion is supported by both the distinct momentum-space pattern—unlike the elastic $(100)$ diffuse scattering—and the negligible magnitude of the estimated leakage relative to the signal in the $[0.2,0.6]$ meV energy window. In Fig. A10(i,j), we also show a similar comparison between the in-plane ($K\leq 0.1$ rlu along $K\bar{K}0]$) powder-averaged scattering intensity as a function of energy transfer $\hbar\omega$ at $T=0.05$ K and the intensity in this $(HHL)$ plane integrated within $\pm 0.1$ rlu around the (111) Bragg peak, scaled to match the elastic peak, for the $E_{i}=0.99$ meV configuration. This comparison again demonstrates that the gapped excitation centered at $\hbar\omega=0.23$ meV is clearly resolved from the elastic line under this configuration.
In Fig. 5(b–d) of the main text, we present constant-energy $\bm{Q}$-maps at $T=0.5$ K and $T=0.05$ K, measured with $E_{i}=3.32$ meV. In these maps, signals from the energy-gain and energy-loss channels have been combined. The main reason we refrain from using only energy-loss signals with $\hbar\omega>0$ and applying the correction factor $(1+\exp(-\hbar\omega/k_{B}T))$ is that it is unclear whether the detailed balance condition, $I(-|\hbar\omega|,-\bm{Q})=\exp(-\hbar\omega/k_{B}T)I(|\hbar\omega|,\bm{Q})$, holds in the spin-frozen phase at $T=0.05$ K. To provide additional information, Fig. A12 presents constant-energy $\bm{Q}$-maps measured with $E_{i}=3.32$ meV, but with the energy-gain and energy-loss channels shown separately. Notably, in the energy transfer range $[0.2,0.6]$ meV [Fig. A12(c,d)], the characteristic “starfish”-shaped magnetic scattering exhibits stronger intensity along the $[111]$ arm compared to the $[001]$ arm. In contrast, the corresponding negative energy range [0.6,0.2][-0.6,-0.2] meV [Fig. A12(i,j)] shows stronger intensity along the $[001]$ arm. Whether this asymmetry reflects non-equilibrium spin dynamics at low temperatures in our $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  sample, or arises from artifacts such as imperfect background subtraction or anisotropic absorption due to the plate-like sample geometry, remains an open question for future investigation.